i must find the sum of this infinite series 9-3+1-1/3+1/9 if there is a sum. please help. i have no idea how to do this
Answer by scott8148(6628)

the ratio (r) of the series is -1/3 ___ this is a converging series because │r│ > 1

(-3) / 9 = -1/3 ___ 1 / (-3) = -1/3 ... etc

the formula for the sum is ___ s = a / (1 - r) ___ where a is the first term

s = 9 / [1 - (-1/3)] ___ s = 9 / (4/3) = 27/4